### v4

@@ 7 Conclusion
Using an algebraic computation system, we have developed a technique for detecting zero-knowledge bugs in arithmetic circuits. This approach is designed to address issues caused by underconstrained or overconstrained arithmetic circuits. Unlike the traditional method of encoding the problem into formulas and utilizing an SMT solver, our approach leverages arithmetic computations, significantly improving efficiency. Additionally, to enhance the verification process, we have incorporated heuristic methods. Notably, our method directly analyzes arithmetic circuits and does not depend on a specific domain-specific language (DSL), making it applicable to a wide range of DSLs that support zk-SNARKs. Our implementation utilizes the open-source algebra computation system, SymPy, in a tool called $\text{AC}^{4}$. We conducted an evaluation involving 176 Circom circuits, and our approach successfully verified approximately 89% of the benchmark sets within a short timeframe.

### v5

@@ 7 Conclusion
Using an algebraic computation system, we have developed a technique for detecting zero-knowledge bugs in arithmetic circuits. This approach is designed to address issues caused by underconstrained or overconstrained arithmetic circuits. Unlike the traditional method of encoding the problem into formulas and utilizing an SMT solver, our approach leverages arithmetic computations, significantly improving efficiency. Additionally, to enhance the verification process, we have incorporated heuristic methods. Notably, our method directly analyzes arithmetic circuits and does not depend on a specific domain-specific language (DSL), making it applicable to a wide range of DSLs that support zk-SNARKs. Our implementation utilizes the open-source algebra computation system, SymPy, in a tool called $\text{AC}^{4}$. We conducted an evaluation involving $176$ Circom circuits and $29$ halo2 circuits, and our approach successfully verified approximately 89% of the benchmark sets within a short timeframe.